Do density, deviation and percent error calculations to 2nd decimal place; average deviation calculation to 3rd decimal place.

Density measurements were conducted on a 23.3oC sample of water which had a theoretical density of 0.997466 g/ml. A volume of 10.00 ml of the water had a mass of 9.92 g. A volume of 15.00 ml of the water had a mass of 15.55 g. A volume of 25.00 ml of the water had a mass of 25.55 g.

(1) Show the calculation of the density of each volume.

(2) Show the calculation of the average density.

(3) Show the calculation of the deviations and the average deviation.

(4) Show the calculation of the percent error based on the theoretical density of 0.997466 g/ml.
1
Expert's answer
2017-03-01T08:54:29-0500
₰₁=m₁/V₁=9.92g/10.00ml= 0.992g/ml=1.00g/ml
₰₂=m₂/V₂=1.0367g/ml= 1.04g/ml
₰₃=m₃/V₃=25.55g/25ml=1.022g/ml= 1.02g/ml
₰middle=0.992+1.367+1.022/3=1.02g/ml
Δ₰₁=0.997466-0.992=0.01g/ml
Δ₰₂=0.997466-1.0367=-0.04g/ml
Δ₰₃=0.997466-1.022=-0.02g/ml
Δ₰middle=0.01-0.04-0.02/3=-0.017g/ml
%error=|0.997466-1.02|*100/0.997466=2.26%